# Anosov representations, strongly convex cocompact groups and weak eigenvalue gaps

### v3

@@ 2.1. Lie theory
For each $1\leq i\leq\frac{d}{2}$ we denote by $P_{i}^{+}$ (resp. $P_{i}^{-}$) the stabilizer of the plane $\langle e_{1},\ldots,e_{i}\rangle$ (resp. $\langle e_{i+1},\ldots,e_{d}\rangle$). The pair of parabolic subgroups $(P_{i}^{+},P_{i}^{-})$ is opposite. An element $g\in\mathsf{GL}_{d}(\mathbb{R})$ is $P_{i}$-proximal if and only if $\ell_{i}(g)>\ell_{i+1}(g)$. In this case $g$ admits a unique attracting fixed point in the space $G/P_{i}^{+}=\mathsf{Gr}_{i}(\mathbb{R}^{d})$.
@@
Let $(X,d)$ be a proper geodesic metric and $x_{0}\in X$ a fixed basepoint. For an isometry $\gamma:X\rightarrow X$ define $|\gamma|_{X}=d(\gamma x_{0},x_{0})$. The translation length and the stable translation length of the isometry $\gamma$ respectively are:
@@
A proper geodesic metric space space $(X,d)$ is called Gromov hyperbolic if there exists $\delta\geqslant 0$ with the following property: for every $x,y,z\in X$
$(x\cdot y)_{x_{0}}\geqslant\min\big{\{}(x\cdot z)_{x_{0}},(z\cdot y)_{x_{0}}\big{\}}-\delta$
and we denote by $\partial_{\infty}X$ the Gromov boundary of $X$.
A finitely generated group $\Gamma$ is called word hyperbolic (or Gromov hyperbolic) if the Cayley graph of $\Gamma$ equipped with the word metric $d_{\Gamma}$ is a Gromov hyperbolic space. In this case, every infinite order element $\gamma\in\Gamma$ has exactly two fixed points $\gamma^{+},\gamma^{-}\in\partial_{\infty}\Gamma$ called the attracting and repelling fixed points of $\gamma$ respectively. For more details on Gromov hyperbolic spaces and their boundaries we refer the reader to [8, Ch. III.H & III.$\Gamma$] and [16].
@@
A function $f:\mathbb{N}\rightarrow(0,\infty)$ is called a Floyd function if it satisfies the following two conditions:
@@
(ii) there exists $\delta\in(0,1)$ such that $\delta f(n+1)\leqslant f(n)\leqslant f(n+1)$ for every $n\in\mathbb{N}$.
Let $\Gamma$ be a finitely generated group. Given a Floyd function $f$ there exists a metric $d_{f}$ on the Cayley graph of $\Gamma$ with respect to $S$ defined as follows (see [19]): for two adjacent vertices $g,h\in\Gamma$ the distance is defined as $d_{f}(g,h)=f(\max\{|g|_{\Gamma},|h|_{\Gamma}\})$. The length of a finite path ${\bf{p}}$ defined by the sequence of adjacent vertices ${\bf{p}}=\{x_{0},x_{1},\ldots,x_{k}\}$ is $L_{f}({\bf{p}})=\sum_{i=0}^{k-1}d_{f}(x_{i},x_{i+1})$. For two arbitrary vertices $g,h\in\Gamma$ their distance is $d_{f}(g,h)=\inf\big{\{}L_{f}({\bf{p}}):{\bf{p}}\ \textup{is a path from}\ g\ \textup{to}\ h\big{\}}$. It is easy to verify that $d_{f}$ defines a metric on $\Gamma$ and let $\overline{\Gamma}$ be the the metric completion of $\Gamma$ with respect to $d_{f}$. Every two points $x,y\in\overline{\Gamma}$ are represented by Cauchy sequences $(\gamma_{n})_{n\in\mathbb{N}},(\delta_{n})_{n\in\mathbb{N}}$ with respect to $d_{f}$ and their distance is $d_{f}(x,y)=\lim_{n}d_{f}(\gamma_{n},\delta_{n})$. The Floyd boundary of $\Gamma$ with respect to $f$ is defined to be the complement $\partial_{f}\Gamma:=\overline{\Gamma}\smallsetminus\Gamma$ equipped with the metric $d_{f}$. The Floyd boundary $\partial_{f}\Gamma$ is called non-trivial if it contains at least three points. For every infinite order element $\gamma\in\Gamma$ the limit $\lim_{n\rightarrow\infty}\gamma^{n}$ exists (see for example [30, Prop. 4]) and is denoted by $\gamma^{+}$.
If $\Gamma$ is a word hyperbolic group, there exists $\epsilon>0$ such that the Floyd boundary of $\Gamma$ with respect to $f(x)=e^{-\epsilon x}$ is the Gromov boundary of $\Gamma$ equipped with a visual metric (see [20]). For more details and properties of the Floyd boundary we refer the reader to [19, 20, 30].
@@
(c) There exist $C,c>0$ such that for every $\hat{m}\in\hat{\Gamma}$, the map $t\mapsto\varphi_{t}(\hat{m})$ is a $(C,c)$-quasi-isometric embedding $(\mathbb{R},d_{\mathbb{E}})\rightarrow(\hat{\Gamma},d_{\hat{\Gamma}})$.
@@
$\big{(}\tau^{+}(\hat{m}),\tau^{-}(\hat{m})\big{)}=\Big{(}\lim_{t\rightarrow\infty}\varphi_{t}(\hat{m}),\ \lim_{t\rightarrow\infty}\varphi_{-t}(\hat{m})\Big{)}$
is well defined, continuous and equivariant with respect to the action of $\Gamma$. For example, if $(M,g)$ is a closed negatively curved Riemannian manifold, a flow space for $\pi_{1}(M)$ satisfying the previous conditions is the unit tangent bundle $T^{1}\widetilde{M}$ equipped with the standard geodesic flow.
Benoist proved that a torsion free, discrete subgroup $\Gamma\subset\mathsf{PGL}_{d}(\mathbb{R})$ acting geometrically on a strictly convex domain $\Omega\subset\mathbb{P}(\mathbb{R}^{d})$ is word hyperbolic (see [5, Thm. 1]). A choice of a flow space for $\Gamma$ is the manifold $T^{1}\Omega$ equipped with the Hilbert geodesic flow.
@@ 2.5. Anosov representations
Let $\rho:\Gamma\rightarrow G$ be a representation and fix $\theta\subset\Delta$ a subset of simple restricted roots of $G$. We denote by $L_{\theta}=P_{\theta}^{+}\cap P_{\theta}^{-}$ the common Levi subgroup. There exists a $G$-equivariant map $G/L_{\theta}\rightarrow G/P_{\theta}^{+}\times G/P_{\theta}^{-}$ mapping the coset $gL_{\theta}$ to the pair $(gP_{\theta}^{-},gP_{\theta}^{-})$. The tangent space of $G/L_{\theta}$ at $(gP_{\theta}^{+},gP_{\theta}^{-})$ splits as the direct sum $\mathsf{T}_{gP_{\theta}^{+}}G/P_{\theta}^{+}\oplus\mathsf{T}_{gP_{\theta}^{-}}G/P_{\theta}^{-}$ and induces a $G$-equivariant splitting of the tangent bundle $\mathsf{T}(G/L_{\theta})=\mathcal{E}\oplus\mathcal{E}^{-}$. We consider the quotient spaces:
@@
where the action of $\gamma\in\Gamma$ on $\mathsf{T}(G/L_{\theta})$ is given by the differential of the left translation by $\rho(\gamma)$ denoted $L_{\rho(\gamma)}:G/L_{\theta}\rightarrow G/L_{\theta}$. Let $\pi:\mathcal{X}_{\rho}\rightarrow\Gamma\backslash\hat{\Gamma}$ and $\pi_{\pm}:\mathcal{E}_{\rho}^{\pm}\rightarrow\mathcal{X}_{\rho}$ be the natural projections. The projections $\pi_{\pm}$ define vector bundles over the space $\mathcal{X}_{\rho}$ where the fiber over the point $[\hat{m},(gP_{\theta}^{+},gP_{\theta}^{-})]_{\Gamma}$ is identified with the vector space $\mathsf{T}_{gP_{\theta}^{\pm}}G/P_{\theta}^{\pm}$. The geodesic flow $\{\varphi_{t}\}_{t\in\mathbb{R}}$ commutes with the action of $\Gamma$ and there exists a lift of the geodesic flow on the quotients $\mathcal{X}_{\rho}$ and $\mathcal{E}_{\rho}^{\pm}$ which we continue to denote by $\{\varphi_{t}\}_{t\in\mathbb{R}}$.
@@
(i) There exists a section $\sigma:\Gamma\backslash\hat{\Gamma}\rightarrow\mathcal{X}_{\rho}$ flat along the flow lines.
(ii) The lift of the geodesic flow $\{\varphi_{t}\}_{t\in\mathbb{R}}$ on the pullback bundle $\sigma_{\ast}\mathcal{E}^{+}$ (resp. $\sigma_{\ast}\mathcal{E}^{-}$) is dilating (resp. contracting).
Two maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$ are called transverse if for any pair of distinct points $(x,y)\in\partial^{(2)}_{\infty}\Gamma$ there exists $h\in G$ such that $(\xi^{-}(x),\xi^{-}(y))=(hP_{\theta}^{+},hP_{\theta}^{-})$. The previous definition is equivalent to the existence of a pair of continuous $\rho$-equivariant transverse maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$ defining the flat section $\sigma:\Gamma\backslash\hat{\Gamma}\rightarrow\mathcal{X}_{\rho}$
$\sigma([\hat{m}]_{\Gamma})=\big{[}\hat{m},(\xi^{+}(\tau^{+}(\hat{m})),\xi^{-}(\tau^{-}(\hat{m})))\big{]}_{\Gamma},$
@@
$\big{|}\big{|}\varphi_{-t}\big{(}X_{v}^{+}\big{)}\big{|}\big{|}_{\varphi_{-t}(x)}\leqslant Ce^{-at}\big{|}\big{|}X_{v}^{+}\big{|}\big{|}_{x}\ \ \ \big{(}\textup{resp.}\ \big{|}\big{|}\varphi_{t}\big{(}X_{v}^{-}\big{)}\big{|}\big{|}_{\varphi_{t}(x)}\leqslant Ce^{-at}\big{|}\big{|}X_{v}^{-}\big{|}\big{|}_{x}\big{)}$
[…]
We recall now some of the key properties of Anosov representations. For more background and for the main properties of Anosov representations see [11, 22, 23, 27, 28, 29, 33]. For a coset $gP_{\theta}^{\pm}$, the stabilizer $\textup{Stab}_{G}(gP_{\theta}^{\pm})$ is the parabolic subgroup $gP_{\theta}^{\pm}g^{-1}$ of $G$. A pair of maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$ are called compatible if for any $x\in\partial_{\infty}\Gamma$ the intersection $\textup{Stab}_{G}(\xi^{+}(x))\cap\textup{Stab}_{G}(\xi^{-}(x))$ is a parabolic subgroup of $G$. We also say that $\xi^{+}$ (resp. $\xi^{-}$) is dynamics preserving if for every infinite order element $\gamma\in\Gamma$, $\rho(\gamma)$ is proximal in $G/P_{\theta}^{+}$ (resp. $G/P_{\theta}^{-}$) and $\xi^{+}(\gamma^{+})$ (resp. $\xi^{-}(\gamma^{+})$) is the attracting fixed point of $\rho(\gamma)$ in $G/P_{\theta}^{+}$ (resp. $G/P_{\theta}^{-}$). We fix an Euclidean norm |||||$\cdot$||| | ⋅ | | on the Cartan subspace $\mathfrak{a}\subset\mathfrak{g}$ and recall that $\mu:G\rightarrow\overline{\mathfrak{a}}^{+}$ denotes the Cartan projection.
@@
(i) There exist $C,c>1$ such that for every $\gamma\in\Gamma$,
$\min_{\alpha\in\theta}\big{\langle}\alpha,\mu(\rho(\gamma))\big{\rangle}\geqslant c^{-1}\big{|}\big{|}\mu(\rho(\gamma))\big{|}\big{|}-c\geq C^{-1}|\gamma|_{\Gamma}-C.$
@@
(ii) $\rho$ admits a pair of compatible, continuous, $\rho$-equivariant maps, dynamics preserving and transverse maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}.$
(iii) The set of $P_{\theta}$-Anosov representations of $\Gamma$ in $G$ is open in $\textup{Hom}(\Gamma,G)$ and the map assigning a $P_{\theta}$-Anosov representation to its Anosov limit maps is continuous.
@@
The following result is the content of [23, Prop. 4.3] and [22, Lem. 3.7] and is used to reduce statements for $P_{\theta}$-Anosov representations to statements for $P_{1}$-Anosov representations.
@@
([22, 23]) Let $G$ a real semisimple Lie group, $\theta\subset\Delta$ a subset of simple restricted roots of $G$. There exists an irreducible $\theta$-compatible representation $\tau:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$, $d=d(G,\theta)$, such that $\tau(P_{\theta}^{+})$ and $\tau(P_{\theta}^{-})$ stabilize the line $[e_{1}]$ and the hyperplane $e_{1}^{\perp}=\langle e_{1},\ldots,e_{d-1}\rangle$ respectively, so that there exist continuous and $\tau$-equivariant embeddings
@@
induced by $\tau$. Moreover, a representation $\rho:\Gamma\rightarrow G$ is $P_{\theta}$-Anosov if and only if $\tau\circ\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ is $P_{1}$-Anosov. In this case, the pair of Anosov limit maps of $\tau\circ\rho$ is $(\iota^{+}\circ\xi^{+},\iota^{-}\circ\xi^{-})$, where $(\xi^{+},\xi^{-})$ is the pair of the limit maps of $\rho$.
@@
The following result was proved by Benoist using a result of Abels–Margulis–Soifer [1] and allows one to control the Cartan projection of a semisimple representation in terms of its Lyapunov projection. We refer the reader to [22, Thm. 4.12] for a proof.
@@
([1] & [3]) Let $G$ be a real reductive Lie group, $\Gamma$ be a discrete group and $\rho_{i}:\Gamma\rightarrow G$, $1\leqslant i\leqslant s$, semisimple representations. Then there exists $C>0$ and a finite subset $F$ of $\Gamma$ such that for every $\gamma\in\Gamma$ there exists $f\in F$ with the property:
$\max_{1\leqslant i\leqslant s}\Big{|}\Big{|}\mu\big{(}\rho_{i}(\gamma)\big{)}-\lambda\big{(}\rho_{i}(\gamma f)\big{)}\Big{|}\Big{|}\leqslant C$
@@
([22, Prop. 1.8]) Let $\Gamma$ be a finitely generated group, $G$ a real semisimple Lie group, $\theta\subset\Delta$ a subset of simple restricted roots of $G$ and $\rho:\Gamma\rightarrow G$ be a representation with semisimplification $\rho^{ss}:\Gamma\rightarrow G$. Then, $\lambda(\rho(\gamma))=\lambda(\rho^{ss}(\gamma))$ for every $\gamma\in\Gamma$ and $\rho$ is $P_{\theta}$-Anosov if and only if $\rho^{ss}$ is $P_{\theta}$-Anosov.

### v4

@@ 2.1. Lie theory
For any integer $1\leq i\leq\frac{d}{2}$ we denote by $P_{i}^{+}$ (resp. $P_{i}^{-}$) the stabilizer of the plane $\langle e_{1},\ldots,e_{i}\rangle$ (resp. $\langle e_{i+1},\ldots,e_{d}\rangle$). The pair of parabolic subgroups $(P_{i}^{+},P_{i}^{-})$ is opposite. An element $g\in\mathsf{GL}_{d}(\mathbb{R})$ is $P_{i}$-proximal if and only if $\ell_{i}(g)>\ell_{i+1}(g)$. In this case $g$ admits a unique attracting fixed point in the flag space $G/P_{i}^{+}=\mathsf{Gr}_{i}(\mathbb{R}^{d})$.
@@
Let $(X,d)$ be a proper geodesic metric and $x_{0}\in X$ a fixed basepoint. For an isometry $\gamma:X\rightarrow X$ define $|\gamma|_{X}:=d(\gamma x_{0},x_{0})$. The translation length and the stable translation length of the isometry $\gamma$ respectively are:
@@
A proper geodesic metric space space $(X,d)$ is called Gromov hyperbolic if there exists $\epsilon\geqslant 0$ with the following property: for every $x,y,z\in X$
$(x\cdot y)_{x_{0}}\geqslant\min\big\{(x\cdot z)_{x_{0}},(z\cdot y)_{x_{0}}\big\}-\epsilon.$
The Gromov boundary of $X$ is denoted by $\partial_{\infty}X$.
[…]
@@
A non-increasing function $f:\mathbb{N}\rightarrow(0,\infty)$ is called a Floyd function if it satisfies the following two conditions:
@@
(ii) there exists $0<\epsilon<1$ such that $\epsilon f(n)\leqslant f(n+1)\leqslant f(n)$ for every $n\in\mathbb{N}$.
Let $\Gamma$ be a finitely generated group. Given a Floyd function $f:\mathbb{N}\rightarrow(0,\infty)$ there exists a metric $d_{f}$ on the Cayley graph of $\Gamma$ with respect to $S$ defined as follows (see [21]): for two adjacent vertices $g,h\in\Gamma$ their distance is defined as $d_{f}(g,h)=f(\max\{|g|_{\Gamma},|h|_{\Gamma}\})$. The length of a finite path $\bf{p}$} defined by the sequence of adjacent vertices ${\bf{p}}=\{x_{0},x_{1},\ldots,x_{k}\}$ is $L_{f}({\bf{p}})=\sum_{i=0}^{k-1}d_{f}(x_{i},x_{i+1})$. For two arbitrary vertices $g,h\in\Gamma$ their distance is $d_{f}(g,h)=\inf\big\{L_{f}({\bf{p}}):{\bf{p}}\ \textup{is a path from}\ g\ \textup{to}\ h\big\}$. It is easy to verify that $d_{f}$ defines a metric on $\Gamma$ and let $\overline{\Gamma}$ be the the metric completion of $\Gamma$ with respect to $d_{f}$. Every two points $x,y\in\overline{\Gamma}$ are represented by Cauchy sequences $(\gamma_{n})_{n\in\mathbb{N}},(\delta_{n})_{n\in\mathbb{N}}$ with respect to $d_{f}$ and their distance is $d_{f}(x,y)=\lim_{n}d_{f}(\gamma_{n},\delta_{n})$. The Floyd boundary of $\Gamma$ with respect to $f$ is defined to be the complement $\partial_{f}\Gamma:=\overline{\Gamma}\smallsetminus\Gamma$ equipped with the metric $d_{f}$. The Floyd boundary $\partial_{f}\Gamma$ is called non-trivial if it contains at least three points. For every infinite order element $\gamma\in\Gamma$ the limit $\lim_{n\rightarrow\infty}\gamma^{n}$ exists (see for example [32, Prop. 4]) and is denoted by $\gamma^{+}$.
If $\Gamma$ is a word hyperbolic group, there exists $\varepsilon>0$ such that the Floyd boundary of $\Gamma$ with respect to $f(x)=e^{-\varepsilon x}$ is the Gromov boundary of $\Gamma$ equipped with a visual metric (see [22]). For more details and properties of the Floyd boundary we refer the reader to [21, 22, 32].
@@
[…]
@@
$)\big(\tau^{+}(\hat{m}),\tau^{-}(\hat{m})\big)=\Big(\lim_{t\rightarrow\infty}\varphi_{t}(\hat{m}),\ \lim_{t\rightarrow\infty}\varphi_{-t}(\hat{m}$$\Big)$
is well-defined, continuous and equivariant with respect to the action of $\Gamma$. For example, if $(M,g)$ is a closed negatively curved Riemannian manifold, a flow space for $\pi_{1}(M)$ satisfying the previous conditions is the unit tangent bundle $T^{1}\widetilde{M}$ equipped with the standard geodesic flow.
Benoist proved that a torsion-free, discrete subgroup $\Gamma\subset\mathsf{PGL}_{d}(\mathbb{R})$ acting geometrically on a strictly convex domain $\Omega\subset\mathbb{P}(\mathbb{R}^{d})$ is word hyperbolic (see [7, Thm. 1]). A choice of a flow space for $\Gamma$ is the manifold $T^{1}\Omega$ equipped with the Hilbert geodesic flow.
@@ 2.5. Anosov representations
Let $\rho:\Gamma\rightarrow G$ be a representation and fix $\theta\subset\Delta$ a subset of simple restricted roots of $G$. We denote by $L_{\theta}=P_{\theta}^{+}\cap P_{\theta}^{-}$ the common Levi subgroup of $P_{\theta}^{+},P_{\theta}^{-}$. There exists a $G$-equivariant embedding $G/L_{\theta}\rightarrow G/P_{\theta}^{+}\times G/P_{\theta}^{-}$ mapping the coset $gL_{\theta}$ to the pair $(gP_{\theta}^{+},gP_{\theta}^{-})$. The tangent space of $G/L_{\theta}$ at $(gP_{\theta}^{+},gP_{\theta}^{-})$ splits as the direct sum $\mathsf{T}_{gP_{\theta}^{+}}G/P_{\theta}^{+}\oplus\mathsf{T}_{gP_{\theta}^{-}}G/P_{\theta}^{-}$ and induces a $G$-equivariant splitting of the tangent bundle $\mathsf{T}(G/L_{\theta})=\mathcal{E}\oplus\mathcal{E}^{-}$. We consider the quotient spaces:
@@
where the action of $\gamma\in\Gamma$ on $\mathsf{T}(G/L_{\theta})$ is given by the differential $dL_{\rho(\gamma)}$ of the left translation by $\rho(\gamma)$, denoted $L_{\rho(\gamma)}:G/L_{\theta}\rightarrow G/L_{\theta}$. Let $\pi:\mathcal{X}_{\rho}\rightarrow\Gamma\backslash\hat{\Gamma}$ and $\pi_{\pm}:\mathcal{E}_{\rho}^{\pm}\rightarrow\mathcal{X}_{\rho}$ be the natural projections. The projections $\pi_{\pm}$ define vector bundles over the space $\mathcal{X}_{\rho}$ where the fiber over the point $\hat{m},(gP_{\theta}^{+},gP_{\theta}^{-})]_{\Gamma}$ is identified with the vector space $\mathsf{T}_{gP_{\theta}^{\pm}}G/P_{\theta}^{\pm}$. The geodesic flow $\{\varphi_{t}\}_{t\in\mathbb{R}}$ commutes with the action of $\Gamma$ and there exists a lift of the geodesic flow on the quotients $\mathcal{X}_{\rho}$ and $\mathcal{E}_{\rho}^{\pm}$ which we continue to denote by $\{\varphi_{t}\}_{t\in\mathbb{R}}$.
@@
  1. (1)
    There exists a section $\sigma:\Gamma\backslash\hat{\Gamma}\rightarrow\mathcal{X}_{\rho}$ flat along the flow lines.
  2. (2)
    The lift of the geodesic flow $\{\varphi_{t}\}_{t\in\mathbb{R}}$ on the pullback bundle $\sigma_{\ast}\mathcal{E}^{+}$ (resp. $\sigma_{\ast}\mathcal{E}^{-}$) is dilating (resp. contracting).
Two maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$ are called transverse if for any pair of distinct points $(x,y)\in\partial^{(2)}_{\infty}\Gamma$ there exists $h\in G$ such that $(\xi^{+}(x),\xi^{-}(y))=(hP_{\theta}^{+},hP_{\theta}^{-})$. The previous definition is equivalent to the existence of a pair of continuous $\rho$-equivariant transverse maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$ defining the flat section $\sigma:\Gamma\backslash\hat{\Gamma}\rightarrow\mathcal{X}_{\rho}$
$\sigma([\hat{m}]_{\Gamma}):=\big[\hat{m},(\xi^{+}(\tau^{+}(\hat{m})),\xi^{-}(\tau^{-}(\hat{m})))\big]_{\Gamma},$
and a continuous equivariant family of norms $(||\cdot||_{x})_{x\in\Gamma\backslash\hat{\Gamma}}$ with the property that there exist $C,a>0$ such that for every $x=[\hat{m}]_{\Gamma}$, $t\geqslant 0$, and $v\in\mathsf{T}_{\xi^{+}(\tau^{+}(\hat{m}))}G/P_{\theta}^{+}$ (resp. $v\in\mathsf{T}_{\xi^{-}(\tau^{-}(\hat{m}))}G/P_{\theta}^{-}$):
@@
where $X_{v}^{+}$ (resp. $X_{v}^{-}$) denotes the copy of the vector $v\in\pi^{-1}_{+}(x)$ $\big(\textup{resp.}\ v\in\pi^{-1}_{-}(x)\big)$.
We recall now some of the key properties of Anosov representations. For more background and for the main properties of Anosov representations see [13, 24, 25, 29, 30, 31, 36]. For a coset $gP_{\theta}^{\pm}$, the stabilizer $\textup{Stab}_{G}(gP_{\theta}^{\pm})$ is the parabolic subgroup $gP_{\theta}^{\pm}g^{-1}$ of $G$. A pair of maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$ are called compatible if for any $x\in\partial_{\infty}\Gamma$ the intersection $\textup{Stab}_{G}(\xi^{+}(x))\cap\textup{Stab}_{G}(\xi^{-}(x))$ is a parabolic subgroup of $G$. We also say that $\xi^{+}$ (resp. $\xi^{-}$) is dynamics preserving if for every infinite order element $\gamma\in\Gamma$, $\rho(\gamma)$ is proximal in $G/P_{\theta}^{+}$ (resp. $G/P_{\theta}^{-}$) and $\xi^{+}(\gamma^{+})$ (resp. $\xi^{-}(\gamma^{+})$) is the attracting fixed point of $\rho(\gamma)$ in $G/P_{\theta}^{+}$ (resp. $G/P_{\theta}^{-}$). We fix an Euclidean norm $||\cdot||$ on the Cartan subspace $\mathfrak{a}\subset\mathfrak{g}$ and recall that $\mu:G\rightarrow\overline{\mathfrak{a}}^{+}$ denotes the Cartan projection.
@@
  1. (i)
    There exist $C,c>1$ such that for every $\gamma\in\Gamma$,
[…]
@@
  2. (ii)
    $\rho$ admits a pair of compatible, continuous, $\rho$-equivariant, dynamics preserving and transverse maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}.$
  3. (iii)
    The set of $P_{\theta}$-Anosov representations of $\Gamma$ in $G$ is open in $\textup{Hom}(\Gamma,G)$ and the map assigning a $P_{\theta}$-Anosov representation to its Anosov limit maps is continuous.
@@
The following result is the content of [25, Prop. 4.3] and [24, Lem. 3.7] and is used to reduce statements for $P_{\theta}$-Anosov representations to statements for $P_{1}$-Anosov representations.
@@
([24, 25]) Let $G$ a real semisimple Lie group, $\theta\subset\Delta$ a subset of simple restricted roots of $G$. There exists an irreducible, $\theta$-compatible representation $\tau_{\theta}:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$, $d=d(G,\theta)$, such that $\tau_{\theta}(P_{\theta}^{+})$ and $\tau_{\theta}(P_{\theta}^{-})$ stabilize the line $e_{1}]$ and the hyperplane $e_{1}^{\perp}=\langle e_{1},\ldots,e_{d-1}\rangle$ respectively, so that there exist continuous and $\tau_{\theta}$-equivariant embeddings
@@
induced by $\tau$. Moreover, a representation $\rho:\Gamma\rightarrow G$ is $P_{\theta}$-Anosov if and only if $\tau_{\theta}\circ\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ is $P_{1}$-Anosov. In this case, the pair of Anosov limit maps of $\tau_{\theta}\circ\rho$ is $(\iota^{+}\circ\xi^{+},\iota^{-}\circ\xi^{-})$, where $(\xi^{+},\xi^{-})$ is the pair of the limit maps of $\rho$.
@@
The following result was proved by Benoist using a result of Abels–Margulis–Soifer [1] and allows one to control the Cartan projection of a semisimple representation in terms of its Lyapunov projection. We refer the reader to [24, Thm. 4.12] for a proof.
@@
([1] & [4]) Let $G$ be a real reductive Lie group, $\Gamma$ be a discrete group and $\big\{\rho_{i}:\Gamma\rightarrow G\big\}_{i=1}^{s}$ semisimple representations. Then there exists $C>0$ and a finite subset $F$ of $\Gamma$ such that for every $\gamma\in\Gamma$ there exists $f\in F$ with the property:
[…]
@@
([24, Prop. 1.8]) Let $\Gamma$ be a finitely generated group, $G$ a real semisimple Lie group, $\theta\subset\Delta$ a subset of simple restricted roots of $G$ and $\rho:\Gamma\rightarrow G$ be a representation with semisimplification $\rho^{ss}:\Gamma\rightarrow G$. Then for every $\gamma\in\Gamma$, $\lambda(\rho(\gamma))=\lambda(\rho^{ss}(\gamma))$ and $\rho$ is $P_{\theta}$-Anosov if and only if $\rho^{ss}$ is $P_{\theta}$-Anosov.